A line passes through (-2,3), (2,5), and (6,k). Find k.
I've been finding the slope of the points, using that information to make a point slope {(y-y)=m(x-x)} equation, and turning that into a y=mx+b equation. However, I am not sure how I should go about solving this one.

Answer by stanbon(57307)

A line passes through (-2,3), (2,5), and (6,k). Find k.
----
slope = (5-3)/(2--2) = 2/4 = 1/2
---
Equation Form:
y = mx = b
5 = (1/2)2 + b
b = 4
----
Equation:
y = (1/2)x+4
---
and (6,k). Find k
--
y = (1/2)6+4 = 7
k = 7
========
Cheers,
Stan H.
=========